# TempoControl: Temporal Attention Guidance for Text-to-Video Models

### v1

@@ 5.5 Conclusions
To test fine-grained temporal modulation of actions, we generate 100 videos using prompts such as: “A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,” where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote subject, verb, adverb, and time reference, respectively. Example: “A child jumping joyfully at the third second.” The verb is controlled via a binary temporal mask (1 only at $\langle t\rangle$), while the subject is always present.
Multi-Object Accuracy.
We benchmark TempoControl on the VBench Multiple Objects benchmark, which evaluates the generation of all the mentioned objects. Each video is sampled at 16 evenly spaced frames. A frame is successful if both target objects are detected. The final score is the average per-frame success rate across all videos. We use all-one control signals for both objects and optimize only the first 4 denoising steps, skipping the Pearson-based stopping criterion.
.9 Human Evaluation Protocol
To assess temporal alignment and visual quality, we conducted a human preference study using side-by-side video comparisons. For each prompt, two videos were presented, one generated by our method and one by the text baseline, randomly assigned as Video A or B to avoid bias.
@@
Annotators were asked to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)? and (2) Which video looks more visually appealing? A “Same (Use sparingly!)” option was available if no difference was perceived.
Example screenshots of the evaluation interface are shown in Figures 6 and 7.
@@
.10 Multi-Object Evaluation
We report extended results on multi-object text-to-video generation using the VC-Bench benchmark. In this setting, prompts reference multiple entities with temporal conditions (e.g., “A dog in the beginning, then a cat”). We compare our method against a text baseline.
Table 4 summarizes results. Our method achieves substantial gains in object-centric metrics. Specifically, we improve multi-object accuracy from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating improved temporal grounding and visibility of both entities. We also observe improvements across all consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
The dynamic degree shows a slight reduction (from 30.5% to 18.8%), which reflects a known tradeoff: our method promotes temporal consistency, often leading to more stable object appearance throughout the video. As a result, motion may be reduced.
However, we note a limitation of this metric: VC-Bench implicitly favors videos where objects are present in all frames. A more suitable benchmark might instead reward appearances rather than requiring full-video presence.
@@
.11 Ablations
Table 5 presents an ablation study in the two-object setting, evaluating the impact of removing the energy term, entropy term, and spatial penalty.
Omitting the energy term results in a modest decrease in temporal accuracy (from 55% to 52.56%). In contrast, removing the entropy term has a more significant effect: temporal accuracy falls to 48.96%, and the temporal presence accuracy drops sharply from 43.54% to 34.76%. Subject Consistency, Aesthetic Quality, and Imaging Quality also decline. Without entropy regularization, attention becomes overly diffuse, often blending the two objects.
Qualitative analysis shows that without the entropy term, the model frequently replaces the static object with the dynamic one (see Section .8 for details on the static and temporally conditioned objects, and Figure 8 for qualitative results).
Table 5: Ablation results for the two-object setting.
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Aesthetic Quality Imaging Quality Text baseline 37.50% 45.85% 29.15% 92.06% 95.32% 98.92% 62.20% 53.26% 68.56% Ours 55.00% 66.46% 43.54% 92.21% 95.43% 99.19% 41.46% 56.55% 70.66% No energy 52.56% 58.90% 46.22% 92.31% 95.45% 99.34% 46.34% 56.12% 71.15% No entropy 48.96% 63.17% 34.76% 89.99% 94.39% 98.92% 53.66% 54.66% 68.95% No spatial 52.01% 60.85% 43.17% 92.76% 95.67% 99.23% 48.78% 56.75% 70.25%
Figure 8: Ablation study. Left: without entropy regularization, the motorcycle disappears. Right: the bear is incorrectly replaced by a zebra.
.12 Impact of Temporal Phrasing
@@
We analyze the effect of using explicit temporal phrasing (e.g., “in the third second,” “in the second part of the video”) in the prompt for the one-object setup. Table 6 compares Wan 2.1 and our method, each evaluated with and without such phrasing.
For Wan 2.1, omitting temporal cues slightly improves temporal accuracy (63.94% to 65.18%) and yields higher aesthetic (45.98% to 48.78%) and imaging quality (53.76% to 59.99%), suggesting that the model struggles to interpret or apply explicit timing effectively.
Our method benefits even more: removing temporal phrases results in the highest temporal accuracy (81.12%) along with gains in aesthetic (48.16%) and imaging quality (56.32%). We hypothesize that Wan 2.1 is not trained to handle precise temporal expressions, and current attention mechanisms may not effectively capture their semantics. As a result, such phrasing introduces ambiguity. In contrast, concise prompts without explicit timing allow our method to align more reliably with the intended temporal structure.
For fairness, all main paper results use prompts with temporal phrasing to match the baseline. However, these findings suggest that our method achieves even stronger performance when using simpler prompts without explicit timing constraints.
@@
Method Temp. Acc. T. Abs. Acc. T. Pres. Acc. Subj. Cons. Backg. Cons. Motion Smooth. Dyn. Deg. Aesthetic Imaging Wan with time 63.94% 67.38% 60.50% 84.66% 92.30% 99.16% 47.50% 45.98% 53.76% Wan without time 65.18% 65.12% 65.25% 85.29% 92.44% 99.22% 45.00% 48.78% 59.99% Ours with time 81.00% 88.75% 72.91% 80.60% 90.33% 99.05% 38.75% 44.37% 52.99% Ours without time 81.12% 84.75% 77.50% 81.29% 90.72% 99.06% 35.00% 48.16% 56.32%
.13 Additional Qualitative Examples
@@
apple 4sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
bear 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
bench 2sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
bus 2sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
clock 5sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
elephant 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
surfboard 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
tie 3sec
TextOurs
umbrella 3sec
@@
TextOurs
skateboard 3sec
@@
a bird and a cat
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
a boat and an airplane
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
a book and a clock
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
a carrot and an umbrella
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a chair and a couch
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a giraffe and a bird
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a horse and a sheep
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
a teddy bear and a frisbee
TextOurs
a train and a boat
@@
TextOurs
an umbrella and a handbag
TextOurs
a bear and a zebra
@@
TextOurs
a cow and an elephant
Refer to caption Refer to caption Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
TextOurs
a cat and a dog
Figure 10: Two objects Text vs Ours.
TextOurs
chimpanze clapping 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
horse shaking its mane 4sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
lion roaring 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
magician pulling rabbit 5sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
man lifting weights 3sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
rabbit thumping 4second
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
snake slithering 1sec
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
TextOurs
tiger leaping 1sec
TextOurs
child jumping 1sec
@@
TextOurs
wolf howling 2sec
Figure 11: Movement Text vs Ours.

### v2

@@ 5.4 Conclusions
To evaluate fine-grained temporal modulation of actions, we generate 100 videos using prompts of the form: ‘A video of $\langle s\rangle$ $\langle v\rangle$ $\langle a\rangle$ with a strong movement at the $\langle t\rangle$ second,’ where $\langle s\rangle$, $\langle v\rangle$, $\langle a\rangle$, and $\langle t\rangle$ denote the subject, verb, adverb, and time reference, respectively.
The full list of subject–verb–adverb triplets is given below:
  1. 1.
    a cat pouncing silently
  2. 2.
    a dancer twirling elegantly
  3. 3.
    a bird flapping frantically
  4. 4.
    a lion roaring loudly
  5. 5.
    a wave crashing violently
  6. 6.
    a ballerina leaping lightly
  7. 7.
    a helicopter spinning noisily
  8. 8.
    a snake slithering smoothly
  9. 9.
    a bear growling deeply
  10. 10.
    a man running fast
  11. 11.
    a woman laughing brightly
  12. 12.
    a child crawling slowly
  13. 13.
    a monkey swinging playfully
  14. 14.
    a firework exploding suddenly
  15. 15.
    a wolf howling hauntingly
  16. 16.
    a dolphin diving swiftly
  17. 17.
    a plane ascending steadily
  18. 18.
    a kangaroo hopping energetically
  19. 19.
    a fox darting quickly
  20. 20.
    a shadow moving mysteriously
  21. 21.
    a squirrel scampering nervously
  22. 22.
    a truck crashing loudly
  23. 23.
    a child screaming unexpectedly
  24. 24.
    a penguin waddling adorably
  25. 25.
    a frog croaking rhythmically
  26. 26.
    a leaf drifting slowly
  27. 27.
    a shark circling menacingly
  28. 28.
    a comet streaking brightly
  29. 29.
    a dancer collapsing dramatically
  30. 30.
    a boxer punching fiercely
  31. 31.
    a tree swaying gently
  32. 32.
    a man lifting weights powerfully
  33. 33.
    a windmill rotating steadily
  34. 34.
    a girl skipping cheerfully
  35. 35.
    a car drifting dangerously
  36. 36.
    a train accelerating fast
  37. 37.
    a snake striking quickly
  38. 38.
    a fire burning intensely
  39. 39.
    a glacier cracking slowly
  40. 40.
    a bee buzzing constantly
  41. 41.
    a deer sprinting fearfully
  42. 42.
    a volcano erupting violently
  43. 43.
    a runner collapsing from exhaustion
  44. 44.
    a gymnast flipping smoothly
  45. 45.
    a rocket launching thunderously
  46. 46.
    a drummer hitting rapidly
  47. 47.
    a magician vanishing mysteriously
  48. 48.
    a spider crawling delicately
  49. 49.
    a tiger pacing restlessly
  50. 50.
    a surfer balancing skillfully
  51. 51.
    a leopard growling softly
  52. 52.
    a swimmer diving gracefully
  53. 53.
    a baby clapping happily
  54. 54.
    a mime gesturing expressively
  55. 55.
    a crow cawing sharply
  56. 56.
    a goat headbutting suddenly
  57. 57.
    a girl blowing bubbles gently
  58. 58.
    a chef chopping rapidly
  59. 59.
    a horse shaking its mane proudly
  60. 60.
    a robot malfunctioning erratically
  61. 61.
    a meteor falling fast
  62. 62.
    a snake coiling tightly
  63. 63.
    a rabbit thumping nervously
  64. 64.
    a man collapsing dramatically
  65. 65.
    a car braking suddenly
  66. 66.
    a hawk diving precisely
  67. 67.
    a wolf snarling viciously
  68. 68.
    a plane landing smoothly
  69. 69.
    a firetruck speeding urgently
  70. 70.
    a chimpanzee clapping playfully
  71. 71.
    a dancer stomping rhythmically
  72. 72.
    a snail inching slowly
  73. 73.
    a lioness crouching quietly
  74. 74.
    a storm cloud rolling ominously
  75. 75.
    a bee landing precisely
  76. 76.
    a magician pulling a rabbit suddenly
  77. 77.
    a glacier melting steadily
  78. 78.
    a jellyfish floating gracefully
  79. 79.
    a tree falling loudly
  80. 80.
    a drummer drumming fiercely
  81. 81.
    a torch blazing brightly
  82. 82.
    a woman spinning dramatically
  83. 83.
    a child running barefoot
  84. 84.
    a fox sniffing cautiously
  85. 85.
    a man throwing a ball forcefully
  86. 86.
    a bear shaking off water heavily
  87. 87.
    a lion walking regally
  88. 88.
    a horse rearing suddenly
  89. 89.
    a falcon gliding silently
  90. 90.
    an elephant raising its trunk highly
  91. 91.
    a dog barking forcefully
  92. 92.
    a man sneezing powerfully
  93. 93.
    a child jumping joyfully
  94. 94.
    a tiger leaping fiercely
  95. 95.
    a woman spinning gracefully
  96. 96.
    a robot marching stiffly
  97. 97.
    a horse galloping wildly
  98. 98.
    an eagle soaring majestically
  99. 99.
    a flame flickering rapidly
  100. 100.
    a skateboarder flipping expertly
The verb is modulated using a binary temporal mask (set to 1 only at time $\langle t\rangle$), while the subject remains active throughout the video.
.8 Human Evaluation Protocol
To assess temporal alignment and visual quality, we conducted a human preference study using side-by-side video comparisons. All annotators were graduate students conducting research in computer vision or closely related areas.
For each prompt, two videos were shown: one generated by our method and one by the text-only baseline, with their order (Video A or Video B) randomized to avoid bias.
@@
Annotators were instructed to read the prompt and answer two questions: (1) Which video is more temporally accurate (i.e., does the object appear at the correct time and remain visible)? (2) Which video is more visually appealing? A “Same (use sparingly)” option was available when no meaningful difference was perceived.
After the experiments, several annotators noted that judging temporal control could be challenging when the difference in timing was only 1–2 seconds. They reported that knowing the exact expected trigger time from the control mask would have helped them assess temporal accuracy more reliably. While we chose not to repeat the study, we include this observation as guidance for future temporal evaluation protocols.
Example screenshots of the evaluation interface are shown in Figures 7 and 8.
@@
.9 Multi-Object Evaluation
We report extended results on multi-object text-to-video generation using the VBench benchmark. In this setting, prompts reference multiple entities (e.g., ‘A dog and a cat’), and the evaluation measures whether all specified objects appear consistently throughout the video. Each video is sampled at 16 evenly spaced frames, and a frame is counted as successful if both target objects are detected; the final score is the average per-frame success rate. In the context of our method, this setup corresponds to applying a constant mask of one for both objects across all frames.
For completeness, we briefly summarize the VBench metrics reported in Table 5:
  • Multiple-Objects Accuracy (GriT / YOLO). Measures whether all referenced objects appear correctly in each sampled frame, using open-vocabulary detectors (GriT) or standard detectors (YOLO). Higher values indicate more reliable object visibility and grounding.
  • Subject Consistency. Quantifies how consistently the appearance of the main subject is preserved across frames. It is computed using feature similarity both between consecutive frames and relative to the first frame of the video; higher values indicate a more stable and coherent subject identity.
  • Background Consistency. Measures the stability of the background across the video. High scores indicate coherent scene structure without abrupt changes.
  • Dynamic Degree. Captures the amount of motion or temporal change in the video. The metric is binary at the video level: a video is marked as dynamic if its average optical flow surpasses a predefined threshold. Higher values indicate more motion, while lower values reflect greater temporal stability.
Table 5 summarizes the results. Our method achieves substantial gains in object-centric metrics: multi-object accuracy improves from 74.1% to 76.4% (GriT) and from 61.5% to 65.7% (YOLO), demonstrating better grounding and more reliable visibility of both entities. We also observe improvements across consistency metrics, including subject consistency (from 97.2% to 97.8%) and background consistency (from 97.6% to 98.1%).
The Dynamic Degree decreases from 30.5% to 18.8%, reflecting a trade-off: our method promotes temporal stability, often leading to more consistent object appearance across the video, at the cost of reduced motion. This highlights a limitation of this evaluation protocol: VBench counts an object as correct only when it is present in every sampled frame; hence, the benchmark implicitly favors unnatural videos in which objects persist throughout the sequence.
@@
.10 Ablations
We conduct an ablation study to analyze the influence of the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$ on temporal control and visual quality. The results reveal clear and complementary effects of the two loss components. Increasing the magnitude term $\lambda_{1}$ reduces Imaging Quality while improving T. Absence Accuracy. This suggests that stronger magnitude guidance more strictly enforces the temporal mask, but may also reduce overall visual fidelity.
In contrast, increasing the entropy term $\lambda_{2}$ improves Imaging Quality and raises T. Presence Accuracy, but it decreases T. Absence Accuracy, indicating that stronger entropy regularization encourages more persistent object appearance.
Overall, our chosen hyperparameters ($\lambda_{1}=0.3$, $\lambda_{2}=10$) provide a balanced tradeoff between these competing effects, achieving strong Temporal Accuracy while maintaining high Imaging Quality.
Table 6: Ablation results for the one object setting. All variants use the same temporal Pearson term; we ablate the magnitude weight $\lambda_{1}$ and the entropy weight $\lambda_{2}$. Best scores per column are in bold.
(a) Effect of magnitude weight $\lambda_{1}$ with fixed entropy weight $\lambda_{2}=10$.
Method
Temporal
Accuracy
T. Absence
Accuracy
T. Presence
Accuracy
Imaging
Quality
Text baseline 63.94% 67.38% 60.50% 53.76%
Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
$\lambda_{1}{=}0.1,\lambda_{2}{=}10$ 79.50% 80.62% 78.38% 56.37%
$\lambda_{1}{=}0.65,\lambda_{2}{=}10$ 81.06% 86.50% 75.62% 53.94%
$\lambda_{1}{=}1.0,\lambda_{2}{=}10$ 83.12% 89.00% 77.25% 53.74%
(b) Effect of entropy weight $\lambda_{2}$ with fixed magnitude weight $\lambda_{1}=0.3$.
Method
Temporal
Accuracy
T. Absence
Accuracy
T. Presence
Accuracy
Imaging
Quality
Text baseline 63.94% 67.38% 60.50% 53.76%
Ours ($\lambda_{1}{=}0.3,\lambda_{2}{=}10$) 82.50% 83.25% 81.75% 56.51%
$\lambda_{1}{=}0.3,\lambda_{2}{=}1$ 80.38% 91.25% 69.50% 52.01%
$\lambda_{1}{=}0.3,\lambda_{2}{=}5$ 81.88% 88.50% 75.25% 54.04%
$\lambda_{1}{=}0.3,\lambda_{2}{=}15$ 81.56% 81.62% 81.50% 56.31%
Figure 9 illustrates ablation results for the entropy term, revealing the limitation of using the Pearson correlation term alone without entropy regularization. While Pearson-based optimization successfully forces the object to appear at the desired time, it often alters the object’s semantics, leading to noticeable inconsistencies. For example, in the “Cake” row, the Pearson-only variant transforms the cake into a noticeably different shape and texture; in the “Sandwich” row, the sandwich becomes distorted across frames; and in the “Book” row, the open book becomes a closed stack with altered proportions. Similar semantic drift is observed in the “Keyboard,” “Oven,” and “Stop sign” examples.
These results highlight that Pearson correlation alone can enforce temporal timing but does not sufficiently constrain the appearance space, resulting in unwanted semantic changes. The entropy regularization helps mitigate those artifacts.
Figure 9: Comparison between Wan2.1 and Pearson-only optimization for single-object temporal control. Pearson correlation alone often satisfies the temporal constraint but alters object semantics, producing distorted or inconsistent appearances (e.g., the cake becoming misshaped, the sandwich warping, the book changing structure, or the keyboard turning into a laptop). Adding entropy regularization prevents these failures by stabilizing the object’s identity across time.
.11 Limitations of Explicit Temporal Cues in State-of-the-Art Text-to-Video Models
@@
We analyze the effect of using explicit temporal phrasing (e.g., ‘in the third second’, ‘in the second part of the video’) in the prompt for the one-object setup. Table 7 compares Wan 2.1, Wan 2.2, and our method, each evaluated with and without such phrasing.
For Wan 2.1, explicit temporal cues do not improve temporal grounding. In fact, removing them increases Temporal Accuracy (63.94% to 65.18%) and leads to a clear gain in Imaging Quality (53.76% to 59.99%), indicating that the model struggles to interpret precise timing expressions and performs better without them.
Wan 2.2 shows a more nuanced pattern: it handles temporal phrasing better than Wan 2.1, likely due to improved data or training techniques. However, the drop in quality when explicit timing is introduced is still significant. Imaging Quality decreases sharply from 57.78% (without time) to 48.13% (with time), showing that temporal phrasing introduces confusion even for stronger models. Presence Accuracy also falls (79.50% to 68%), further highlighting that explicit timing signals continue to degrade performance.
Our method remains the most robust and accurate across settings. Across all conditions, our method provides the most reliable and precise temporal control, regardless of prompt style. Interestingly, although our approach builds on Wan 2.1 as the base model, explicit temporal phrasing does increase the temporal accuracy. This suggests that our optimization procedure may curate the base model’s timing capabilities.
@@
Method Temp. Acc. T. Abs. Acc. T. Pres. Acc. Imaging Quality
Wan2.1 with time 63.94% 67.38% 60.50% 53.76%
Wan2.1 without time 65.18% 65.12% 65.25% 59.99%
Wan2.2 with time 78.02% 87.42% 68.00% 48.13%
Wan2.2 without time 75.24% 71.25% 79.50% 57.78%
Ours with time 82.50% 83.25% 81.75% 56.51%
Ours without time 78.75% 82.25% 75.25% 57.14%
.12 Additional Qualitative Examples
@@
a baseball glove and a skateboard
a bicycle and a car
a bicycle and an airplane
a bird and a cat
a boat and an airplane
a book and a clock
a bowl and a remote
a bus and a traffic light
a cake and a vase
a car and a motorcycle
a cat and a dog
a cell phone and a book
a donut and a suitcase
Figure 10: Two Objects Text vs Ours.
@@
a giraffe and a bird
a handbag and a tie
a horse and a sheep
a laptop and a remote
a motorcycle and a boat
a motorcycle and a bus
a person and a toothbrush
a pizza and a tie
a potted plant and a tv
a spoon and a laptop
a tennis racket and a bottle
a toaster and a teddy bear
a truck and a bicycle
a zebra and a giraffe
Figure 11: Two Objects Text vs Ours.
@@
airplane 2nd second
@@
bus 2nd second
cell phone last second
clock last second
cow 2nd second
dining table last second
dog last second
donut 4th second
fork 4th second
@@
hair drier last second
handbag 3rd second
kite 3rd second
knife 4th second
scissors last second
sheep 2nd second
skateboard 3rd second
snowboard 3rd second
sports ball 3rd second
suitacase 3rd second
toilet last second
truck 2nd second
umbrella 3rd second
Figure 13: One Object Text vs Ours.
@@
bee buzzing 3rd second
bird flapping 2nd second
child crawling 2nd second
child jumping 1st second
child screaming 2nd second
chimpanze clapping 1st second
crow cawing 4th second
falcon gliding last second
fox darting 2nd second
horse shaking its mane 4th second
jellyfish floating last second
kangaroo hopping 2nd second
leopard growling 4th second
Figure 14: Movement Text vs Ours.
@@
lion roaring 1st second
lion walking last second
magician pulling rabbit last second
penguin waddling 1st second
rabbit thumping 4th second
robot marching 1st second
rocket launching 3rd second
surfer balancing 4th second
swimmer diving 4th second
wolf snarling 4th second
Figure 15: Movement Text vs Ours.
@@
elephant ex1
@@
rooster ex1
@@
rooster ex2
@@
Figure 16: Audio-video Alignment Text vs Ours.